OpenStudy (anonymous):

Crane A can unload a ship in 10 hours and crane B can unload it in 14 hours. How long will it take the two cranes to unload the ship working together? _______________________________________________________________________________ I just need to know how to set this up into an equation.

OpenStudy (whpalmer4):

The key concept here is "find the unit rate". A can unload 1 ship in 10 hours, so its unit rate is 1 ship/10 hours or 0.1 ships/hr. B can unload 1 ship in 14 hours, so its unit rate is 1 ship/14 hours or 0.07142857 ships/hr. A and B working together in perfect harmony ought to be able to unload (1/10 + 1/14) ships per hour.

OpenStudy (whpalmer4):

Here's a similar example. Bob can mow the lawn in 2 hours, Sam can mow the lawn in 3 hours. How long for them to mow it together? Bob's rate = 1 lawn/2 hours or 1/2 Sam's rate = 1 lawn/3 hours or 1/3 Bob and Sam together mow (1/2 + 1/3) lawns per hour. Make a common denominator to add the fractions:\[\frac{1}{2}*\frac{3}{3} + \frac{1}{3}*\frac{2}{2} = \frac{3}{6}+\frac{2}{6} = \frac{5}{6} \text{ lawns/hr}\] \(1\text{ lawn}/ (5/6\text{ lawns/hr}) = 1\text{ lawn} * \frac{6\text{ hr}}{5\text{ lawns}} = \frac{6}{5} \text{ hours}\) to mow the 1 lawn working together.

OpenStudy (anonymous):

so what's the equation I don't really get anything you just said...

OpenStudy (amistre64):

this one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I don't really understand what wpalmer was saying

OpenStudy (amistre64):

Crane A can unload a ship in 10 hours crane B can unload it in 14 hours How long will it take the two cranes to unload the ship working together? spose a crane works at an even pace; if it takes 10 hours to complete 1 job, how much of the job is completed in 1 hour? in otherwords, what is the hourly rate?

OpenStudy (whpalmer4):

Every hour, crane A unloads 1/10 of the ship. Every hour, crane B unloads 1/14 of the ship. Together, every hour they unload 1/10 + 1/14 of the ship. Add 1/10 + 1/14. Now, divide 1 by that number to give you the number of hours it takes to unload the ship.

OpenStudy (anonymous):

so (1/10 + 1/14) ______________ ? 1

OpenStudy (whpalmer4):

To divide by a fraction, invert it and multiply instead. For example, 1 / (2/3) can be found by multiplying 1 by (3/2).

OpenStudy (whpalmer4):

Close, you want: \[\frac{ 1} {\frac{1}{10} + \frac{1}{14}}\]

OpenStudy (anonymous):

oh but wouldn't that just equal 1/10 + 1/14

OpenStudy (whpalmer4):

Certainly not!

OpenStudy (whpalmer4):

1/10 = 0.1. 1/14 = 0.0714285 Add the two together and you get about 0.17. 1/0.17 does not equal 0.17...

OpenStudy (whpalmer4):

For the purposes of getting a handle on this, say both cranes work at the same speed as crane A. Crane A does 1/10 of the ship each hour. Crane B does 1/10 of the ship each hour. Together, they do 1/10 + 1/10 = 2/10 of the ship per hour. How long would they take to do the entire ship?

OpenStudy (whpalmer4):